OpenStudy (anonymous):

How do i graph a circle when i have the center of 4,3 and radius=6. What are the intercepts

OpenStudy (anonymous):

\[(x - h)^{2} +(y-k)^{2} = r ^{2}\] this is the general equation of a circle centered at (h, k) and radius r. does this information help?

OpenStudy (anonymous):

to graph, you could either just plot points of that equation -- or start at the center (4, 3), and then go 6 units in any direction and draw a circle.

OpenStudy (anonymous):

not really i have the circle graphed but i dont know how to find all the four intersections

OpenStudy (anonymous):

like x and y intercepts? to find x-intercepts, let y = 0 and solve for x. for y-intercepts, let x = 0 and solve for y. how many solutions you get depend on how many times the circle intersects the axes

OpenStudy (anonymous):

what is it i dont really know how to put it in a calculator

OpenStudy (anonymous):

what kind of calculator do you use

OpenStudy (anonymous):

ti 84

OpenStudy (anonymous):

ok do you agree that the equation of this circle is \[(x - 4)^{2} + (y - 3)^{2} = 36\] ?

OpenStudy (anonymous):

i think so because i started off with x^2+y^2-8x-6y-11=0

OpenStudy (anonymous):

that is correct. to enter it into your calculator you will have to solve for y. \[(y - 3)^{2} = 36 - (x - 4)^{2}\] \[y = 3 \pm \sqrt{36 - (x - 4)^{2}}\] so you will have to enter two equations into your calculator

OpenStudy (anonymous):

and what do i enter?

OpenStudy (anonymous):

y = 3 + sqrt(36 -(x - 4)^2) and y = 3 - sqrt(36 - (x - 4)^2)

OpenStudy (anonymous):

each equation will graph 1/2 of the circle. both of them will generate a complete circle. the reason only half can be generated at a time is because a circle is not a function (it fails the vertical line test) but when you cut the circle in half horizontally, you get two functions that, combined, generate the whole circle

OpenStudy (anonymous):

Hmm what is the 4 intersects